OpenStudy (anonymous):

A subway train starts from rest at a station and accelerates at a rate of 1.55 m/s^2 for 13.9 s. It runs at constant speed for 70.3 s and slows down at a rate of 3.51 m/s^2 until it stops at the next station. Find the total distance covered!

OpenStudy (amistre64):

accelerates at a rate of 1.55 m/s^2 for 13.9 s. a = 1.55 velocity=\(\int\)a = 1.55t position=\(\int\)v = .725 t^2 13.9 13.9 ------- 1251 417 139 -------- 193.21 .725 --------- 96605 38642 135247 ----------- 140.07725 meters so far, if i did it right the speed hes going at this moment is: v = 1.55 13.9 ----- 1395 465 155 -------- 21.545 meters per sec. It runs at constant speed for 70.3 s 21.545 70.3 ------- 6435 1508150 ---------- 1508.7935 meters more +140.07725 ------------ 1648.87075 meters total ... if im doing it right and slows down at a rate of 3.51 m/s^2 until it stops at the next station. -3.51t + 21.545 = 0 when t=21.545/3.51, gonna try a calculator on this one :) in 6.14 seconds or so -1.755t^2 should give us the final stretch, and of course a calulator ... 1.755(6.14)^2 = 66.1628 meters more 1648.87075 66.162798 ------------- 1715.033548 meters maybe?

OpenStudy (amistre64):

anyway to check it?

OpenStudy (anonymous):

You're 100% correct, and but I ended up with 149.738 for the first leg of the trip. Not sure how, going to have to go back and check my work!

OpenStudy (amistre64):

thnx, i think the math tho is typoed in the middle1654.69075 is what the calulator tells me.. so im a little off i think

OpenStudy (anonymous):

Thank you, I am going to look over your work and see where I went wrong! Wonderful job!